OpenStudy (anonymous):

substitutions help again! Did this 4 times different answer every time 3x+2y=23 1/2x-4=y

OpenStudy (amistre64):

what does "y" equal in relation to "x" ?

OpenStudy (amistre64):

"(1/2)x -4" = "y" right?

OpenStudy (anonymous):

I can see that, I used this as a substitute for y and keep getting different answers.

OpenStudy (anonymous):

\[3x+2y=23\] put in Y \[3x+2(1/2x-4)\] distribute the 2*1/2x and 2*-4 \[3x+x-8=23\] solve

OpenStudy (amistre64):

yep, then its the adding thats giving you troubles :) happens to me alot still on my calculus tests